# Matrix Rosenthal and Concentration Inequalities for Markov Chains with Applications in Statistical Learning

### v1

@@ Appendix B Omitted Proof for the Matrix Rosenthal Inequality for Markov Chains
    $=$ $\sup_{z\in{\mathcal{Z}}}\sum_{k=0}^{\infty}\left\|{{\mathrm{Q}}^{k}(z,\cdot)({\bm{F}})-\pi({\bm{F}})}\right\|$
    $\leq$ $\sup_{z\in{\mathcal{Z}}}\sum_{k=0}^{\infty}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}\left\|{{\mathrm{Q}}^{k}(z,\cdot)-\pi}\right\|_{\textup{TV}}$
    $\leq$ $\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}\sum_{k=0}^{\infty}\sup_{z\in{\mathcal{Z}}}\left\|{{\mathrm{Q}}^{k}(z,\cdot)-\pi}\right\|_{\textup{TV}}$
@@
    $=$ $\sum_{k=0}^{\infty}\left\|{{\mathrm{Q}}^{k}(z,\cdot)({\bm{F}}{\bm{u}})-\pi({\bm{F}}{\bm{u}})}\right\|$
    $\leq$ $\sum_{k=0}^{\infty}\left\|{\left\|{{\bm{F}}{\bm{u}}}\right\|}\right\|_{\infty}\left\|{{\mathrm{Q}}^{k}(z,\cdot)-\pi}\right\|_{\textup{TV}}$
@@
    $\left\|{\xi{\mathrm{Q}}^{k}-\pi}\right\|_{V^{\alpha}}\leq 2^{1-\alpha}\varkappa^{\alpha}\left({\pi(V)+\xi(V)}\right)^{\alpha}(1/4)^{{\lfloor k/t_{\operatorname{mix}}\rfloor}\alpha}.$
@@
    $\leq$ $\sum_{k=0}^{\infty}\left\|{{\mathrm{Q}}^{k}(z,\cdot)({\bm{F}})-\pi({\bm{F}})}\right\|$
    $\leq$ $\sum_{k=0}^{\infty}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{\alpha}}\left\|{{\mathrm{Q}}^{k}(z,\cdot)-\pi}\right\|_{V^{\alpha}}$
    $\leq$ $2^{1-\alpha}\varkappa^{\alpha}\left({\pi(V)+V(z)}\right)^{\alpha}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{\alpha}}\sum_{k=0}^{\infty}(1/4)^{{\lfloor k/t_{\operatorname{mix}}\rfloor}\alpha}$
    $\leq$ $\left[{\varkappa^{\alpha}2^{3-\alpha}/(3\alpha)}\right](\pi(V)+V(z))^{\alpha}t_{\operatorname{mix}}\left\|{\|{\bm{F}}\|}\right\|_{V^{\alpha}},$
@@
    $\left\|{\left\|{{\bm{S}}_{n}}\right\|}\right\|_{\pi,p}\leq\left\|{\left\|{\sum_{k=0}^{\lfloor{n}/{t_{\operatorname{mix}}}\rfloor t_{\operatorname{mix}}-1}{{\bm{F}}}(Z_{k})}\right\|}\right\|_{\pi,p}+t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}.$
@@
    $\left\|{\left\|{\sum_{k=0}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor t_{\operatorname{mix}}-1}{{\bm{F}}}(Z_{k})}\right\|}\right\|_{\pi,p}$ $=\left\|{\left\|{\sum_{r=0}^{t_{\operatorname{mix}}-1}\sum_{k=0}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor-1}\left({{{\bm{G}}_{t_{\operatorname{mix}}}}(Z_{kt_{\operatorname{mix}}+r})-\left({{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}}\right)(Z_{kt_{\operatorname{mix}}+r})}\right)}\right\|}\right\|_{\pi,p}$
    $\leq\sum_{r=0}^{t_{\operatorname{mix}}-1}\left\|{\left\|{\sum_{k=0}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor-1}\left({{{\bm{G}}_{t_{\operatorname{mix}}}}(Z_{kt_{\operatorname{mix}}+r})-\left({{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}}\right)(Z_{kt_{\operatorname{mix}}+r})}\right)}\right\|}\right\|_{\pi,p}$
    $\leq\sum_{r=0}^{t_{\operatorname{mix}}-1}\left\|{\left\|{\sum_{k=1}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor}\left({{{\bm{G}}_{t_{\operatorname{mix}}}}(Z_{kt_{\operatorname{mix}}+r})-\left({{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}}\right)(Z_{(k-1)t_{\operatorname{mix}}+r})}\right)}\right\|}\right\|_{\pi,p}+2t_{\operatorname{mix}}\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}.$
@@
    $\left\|{\left\|{\sum_{k=1}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor}\left({{{\bm{G}}_{t_{\operatorname{mix}}}}(Z_{kt_{\operatorname{mix}}+r})-\left({{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}}\right)(Z_{(k-1)t_{\operatorname{mix}}+r})}\right)}\right\|}\right\|_{\pi,p}$
    $\leq$ $C_{R,1}\sqrt{p\vee\log d}\left\|{\left\|{\sum_{k=1}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor}{\mathbb{E}}\left[{\left({{{\bm{G}}_{t_{\operatorname{mix}}}}(Z_{kt_{\operatorname{mix}}+r})-\left({{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}}\right)(Z_{(k-1)t_{\operatorname{mix}}+r})}\right)^{2}\big{|}{{\mathscr{F}}_{r,k-1}}}\right]}\right\|^{1/2}}\right\|_{\pi,p}$
    $+C_{R,2}(p\vee\log d)\left\|{\sup_{k\leq\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor}\|{{\bm{G}}_{t_{\operatorname{mix}}}}(Z_{kt_{\operatorname{mix}}+r})-\left({{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}}\right)(Z_{(k-1)t_{\operatorname{mix}}+r})\|}\right\|_{\pi,p}$
    $\leq$ $C_{R,1}\sqrt{p\vee\log d}\Big{\|}\left({\sum_{k=1}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor}\left({2\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}}\right)^{2}}\right)^{1/2}\Big{\|}_{\pi,p}+C_{R,2}(p\vee\log d)\left\|{2\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}}\right\|_{\pi,p}$
    $=$ $2C_{R,1}\sqrt{(p\vee\log d)\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor}\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}+2C_{R,2}(p\vee\log d)\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}.$
@@
    $\left\|{\left\|{{\bm{S}}_{n}}\right\|}\right\|_{\pi,p}$ $\leq t_{\operatorname{mix}}(2C_{R,1}\sqrt{(p{\vee}\log d)\lfloor{n}/{t_{\operatorname{mix}}}\rfloor}\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}{+}2C_{R,2}(p{\vee}\log d)\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}){+}2t_{\operatorname{mix}}\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}{+}t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}$
    $\leq\frac{16}{3}C_{R,1}\sqrt{(p\vee\log d)nt_{\operatorname{mix}}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}+\left({\frac{16}{3}C_{R,2}(p\vee\log d)+\frac{19}{3}}\right)t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}.$
    Furthermore, if there exists an ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$ for all $z\in{\mathcal{Z}}$, then by Lemma 4.2 (a), it holds that ${\bm{G}}^{2}_{t_{\operatorname{mix}}}(z)\preccurlyeq(64/9){\bm{\Upsilon}}$ for all $z\in{\mathcal{Z}}$. By Theorem 2.1 (b), the above inequality still holds if we replace $d$ with $e\cdot r([(64/9){\bm{\Upsilon}}]\wedge[(1/\lfloor{n}/{t_{\operatorname{mix}}}\rfloor)\|\|\langle{{\bm{M}}^{(t_{\operatorname{mix}})}}\rangle_{\lfloor{n}/{t_{\operatorname{mix}}}\rfloor}\|\|_{\pi,\frac{p}{2}}])$, and double the right-hand side.

### v2

@@ Appendix B Omitted Proof for the Matrix Rosenthal Inequality for Markov Chains
    $=$ $\sup_{z\in{\mathcal{Z}}}\sum_{k=0}^{\infty}\left\|{{\mathrm{Q}}^{k}(z,\cdot)({\bm{F}})-\mu({\bm{F}})}\right\|$
    $\leq$ $\sup_{z\in{\mathcal{Z}}}\sum_{k=0}^{\infty}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}\left\|{{\mathrm{Q}}^{k}(z,\cdot)-\mu}\right\|_{\textup{TV}}$
    $\leq$ $\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}\sum_{k=0}^{\infty}\sup_{z\in{\mathcal{Z}}}\left\|{{\mathrm{Q}}^{k}(z,\cdot)-\mu}\right\|_{\textup{TV}}$
@@
    $=$ $\sum_{k=0}^{\infty}\left\|{{\mathrm{Q}}^{k}(z,\cdot)({\bm{F}}{\bm{u}})-\mu({\bm{F}}{\bm{u}})}\right\|$
    $\leq$ $\sum_{k=0}^{\infty}\left\|{\left\|{{\bm{F}}{\bm{u}}}\right\|}\right\|_{\infty}\left\|{{\mathrm{Q}}^{k}(z,\cdot)-\mu}\right\|_{\textup{TV}}$
@@
    $\left\|{\xi{\mathrm{Q}}^{k}-\mu}\right\|_{V^{\alpha}}\leq 2^{1-\alpha}\varkappa^{\alpha}\left({\mu(V)+\xi(V)}\right)^{\alpha}(1/4)^{{\lfloor k/t_{\operatorname{mix}}\rfloor}\alpha}.$
@@
    $\leq$ $\sum_{k=0}^{\infty}\left\|{{\mathrm{Q}}^{k}(z,\cdot)({\bm{F}})-\mu({\bm{F}})}\right\|$
    $\leq$ $\sum_{k=0}^{\infty}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{\alpha}}\left\|{{\mathrm{Q}}^{k}(z,\cdot)-\mu}\right\|_{V^{\alpha}}$
    $\leq$ $2^{1-\alpha}\varkappa^{\alpha}\left({\mu(V)+V(z)}\right)^{\alpha}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{\alpha}}\sum_{k=0}^{\infty}(1/4)^{{\lfloor k/t_{\operatorname{mix}}\rfloor}\alpha}$
    $\leq$ $\left[{\varkappa^{\alpha}2^{3-\alpha}/(3\alpha)}\right](\mu(V)+V(z))^{\alpha}t_{\operatorname{mix}}\left\|{\|{\bm{F}}\|}\right\|_{V^{\alpha}},$
@@
    $\left\|{\left\|{{\bm{S}}_{n}}\right\|}\right\|_{\mu,p}\leq\left\|{\left\|{\sum_{k=0}^{\lfloor{n}/{t_{\operatorname{mix}}}\rfloor t_{\operatorname{mix}}-1}{{\bm{F}}}(Z_{k})}\right\|}\right\|_{\mu,p}+t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}.$
@@
    $\left\|{\left\|{\sum_{k=0}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor t_{\operatorname{mix}}-1}{{\bm{F}}}(Z_{k})}\right\|}\right\|_{\mu,p}$ $=\left\|{\left\|{\sum_{r=0}^{t_{\operatorname{mix}}-1}\sum_{k=0}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor-1}\left({{{\bm{G}}_{t_{\operatorname{mix}}}}(Z_{kt_{\operatorname{mix}}+r})-\left({{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}}\right)(Z_{kt_{\operatorname{mix}}+r})}\right)}\right\|}\right\|_{\mu,p}$
    $\leq\sum_{r=0}^{t_{\operatorname{mix}}-1}\left\|{\left\|{\sum_{k=0}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor-1}\left({{{\bm{G}}_{t_{\operatorname{mix}}}}(Z_{kt_{\operatorname{mix}}+r})-\left({{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}}\right)(Z_{kt_{\operatorname{mix}}+r})}\right)}\right\|}\right\|_{\mu,p}$
    $\leq\sum_{r=0}^{t_{\operatorname{mix}}-1}\left\|{\left\|{\sum_{k=1}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor}\left({{{\bm{G}}_{t_{\operatorname{mix}}}}(Z_{kt_{\operatorname{mix}}+r})-\left({{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}}\right)(Z_{(k-1)t_{\operatorname{mix}}+r})}\right)}\right\|}\right\|_{\mu,p}+2t_{\operatorname{mix}}\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}.$
@@
    $\left\|{\left\|{\sum_{k=1}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor}\left({{{\bm{G}}_{t_{\operatorname{mix}}}}(Z_{kt_{\operatorname{mix}}+r})-\left({{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}}\right)(Z_{(k-1)t_{\operatorname{mix}}+r})}\right)}\right\|}\right\|_{\mu,p}$
    $\leq$ $C_{1}\sqrt{p\vee\log d}\left\|{\left\|{\sum_{k=1}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor}{\mathbb{E}}\left[{\left({{{\bm{G}}_{t_{\operatorname{mix}}}}(Z_{kt_{\operatorname{mix}}+r})-\left({{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}}\right)(Z_{(k-1)t_{\operatorname{mix}}+r})}\right)^{2}\big|{{\mathscr{F}}_{r,k-1}}}\right]}\right\|^{1/2}}\right\|_{\mu,p}$
    $+C_{2}(p\vee\log d)\left\|{\sup_{k\leq\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor}\|{{\bm{G}}_{t_{\operatorname{mix}}}}(Z_{kt_{\operatorname{mix}}+r})-\left({{\mathrm{Q}}^{t_{\operatorname{mix}}}{\bm{G}}_{t_{\operatorname{mix}}}}\right)(Z_{(k-1)t_{\operatorname{mix}}+r})\|}\right\|_{\mu,p}$
    $\leq$ $C_{1}\sqrt{p\vee\log d}\Big\|\left({\sum_{k=1}^{\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor}\left({2\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}}\right)^{2}}\right)^{1/2}\Big\|_{\mu,p}+C_{2}(p\vee\log d)\left\|{2\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}}\right\|_{\mu,p}$
    $=$ $2C_{1}\sqrt{(p\vee\log d)\lfloor\frac{n}{t_{\operatorname{mix}}}\rfloor}\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}+2C_{2}(p\vee\log d)\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}.$
@@
    $\left\|{\left\|{{\bm{S}}_{n}}\right\|}\right\|_{\mu,p}$ $\leq t_{\operatorname{mix}}(2C_{1}\sqrt{(p{\vee}\log d)\lfloor{n}/{t_{\operatorname{mix}}}\rfloor}\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}{+}2C_{2}(p{\vee}\log d)\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}){+}2t_{\operatorname{mix}}\left\|{\left\|{{\bm{G}}_{t_{\operatorname{mix}}}}\right\|}\right\|_{\infty}{+}t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}$
    $\leq\frac{16}{3}C_{1}\sqrt{(p\vee\log d)nt_{\operatorname{mix}}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}+\left({\frac{16}{3}C_{2}(p\vee\log d)+\frac{19}{3}}\right)t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}.$
    Furthermore, if there exists an ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$ for all $z\in{\mathcal{Z}}$, then by Lemma 4.2 (a), it holds that ${\bm{G}}^{2}_{t_{\operatorname{mix}}}(z)\preccurlyeq(64/9){\bm{\Upsilon}}$ for all $z\in{\mathcal{Z}}$. By Theorem 2.1 (b), the above inequality still holds if we replace $d$ with $e\cdot r([(64/9){\bm{\Upsilon}}]\wedge[(1/\lfloor{n}/{t_{\operatorname{mix}}}\rfloor)\|\|\langle{{\bm{M}}^{(t_{\operatorname{mix}})}}\rangle_{\lfloor{n}/{t_{\operatorname{mix}}}\rfloor}\|\|_{\mu,\frac{p}{2}}])$, and double the right-hand side.